What is the pH of a buffer solution that is 0.24M of NH3 and 0.20M of NH4Cl?
1
Expert's answer
2020-12-06T10:35:17-0500

"pKa of NH3 = 4.74\npH = 4.74 + log(NH4Cl\/NH3)\npH = 4.74 + log(0.20\/0.24)\npH = 4.74 + (-0.08)\npH = 4.66"